1. Find the focus and the directx of the graph of each equation. Sketch the graph.
y = 5x^2
2. Find the focus and the directx of the graph of each equation. Sketch the graph.
x = -1/8y^2

Answer by ewatrrr(24785)

pre>
Hi,
y = 5x^2, V(0,0) 5 = 1/4p, p = 1/20. y = -1/20 the directrix (opening Up)
x = (-1/8)y^2 V(0,0) , p = -2,, directrix x = 2 (opening left)